OpenStudy (nepurrta):

Find the missing value to the nearest hundredth. cos ___ = 6/19

OpenStudy (anonymous):

\[\cos^{-1} (\frac{ 6 }{ 19})\]= 71.59151983

OpenStudy (anonymous):

\[\cos^{-1} (\frac{ 6 }{ 19 })\] =71.591 =71.60